Monthly Newsletter

April 2001



Meetings
SORCE is planning a Science Team Meeting July 9, 10, 11 2001 at the Laboratory for Atmospheric and Space Physics, University of Colorado, Boulder, Colorado.

SORCE scientists hope to present papers or attend the following summer 2001 meetings:
  • Workshop on the Evolving Sun and its Influence on Planetary Environments June 2001 Granada, Spain.
  • International Solar Cycle Studies 2001-Solar Variability, Climate and Spaceweather June 2001 Longmont, Colorado
  • IAGA Session on Solar Variability August 2001 Hanoi, Vietnam
SORCE April 2001 Project Summary
The EOS SORCE Mission remains on schedule and within budget. However, virtually all slack time for build, spacecraft integration, and test is dwindling, as are contingency funds. All schedules, manufacturing costs and test facilities are carefully monitored and adjusted on a daily basis to meet the ongoing requirements of the SORCE program. Happily, the extra vigilance and planning has alleviated some situations, and aided others.
On March 28, 2001, Dr. Doug Rabin, new SORCE Deputy Project NASA scientist, met with Dr. Gary Rottman, SORCE PI to discuss the progress of the SORCE program, and Dr. Rottman also met with Dr. Micheal King (NASA EOS Project Scientist) to discuss funding for SORCE for the years 2002 and beyond.
Dr. Myriam Adda, joined the LASP SORCE Mission Team April 23, 2001, as a Research Associate. Dr. Adda will be working with Dr. Jerry Harder on the calibration and characterization of the SIM (Spectral Irradiance Monitor) instrument. Dr. Adda worked on her Ph.D. at the French Space Agency. The objective of her Ph.D. was to design the on-board scientific software for a French satellite, COROT, which deals with stellar seismology and the search for new planets. To obtain this objective, Dr. Adda studied, documented and modeled the perturbations of the instrument, (attitude, stability, CCD sensors) and the instrument environment (thermal variations, diffused light, radiation effects). In addition, Dr. Adda developed the test bench used to test and characterize the CCD sensors for COROT.

SORCE Mission Operations Status
The SORCE Mission Operations Review (MOR) was successfully completed April 23-24, 2001. Eight RFA's (requests for action) resulted from the review.

Two RFA's were for Orbital and dealt with modes/states and the wiring of battery cells, one was directed to NASA regarding the completion of the Project Service Level Agreement, and three RFA's were directed to LASP/SORCE. The three dealt with launch critical assets, launch dispersion analysis, and a possible ICESat/SORCE launch conflict. Realistic launch dates must be obtained to enable smooth launch coordination and mission operation planning among other NASA missions, such as AQUA and ICESat, and mission control operations at LASP, who will control both SORCE and ICESat. Two RFA's were directed to LASP/Orbital and they dealt with the FOG (gyros) review and FDC (failure detection and correction).
The SORCE Mission Operations security plan is accepted, security scans have been planned, and closed IO (input/output) connections should be available in early May, 2001. NASA continues to work on three PSLA issues, the costing of the FDF (flight dynamics facility), the price of voice and data lines and the statement of work for the USN (Universal Space Network). The USN site in Santiago, Chile has been removed from the SORCE LEOP (launch and early orbit phase) and routine operations plan. This will affect operating hours for recieving data from the SORCE satellite. It was stated, that considering the costs for "renting" USN facilities, perhaps SORCE/LASP should consider building a tracking station of their own. It appears that a cost comparison is worth investigation.

TIM Status
As was stated in last month's newsletter TIM saw "first light" March 29, 2001. All TIM "bugs" are reported, tracked and recorded. Everything but the vacuum doors seal, and the vacuum door assembly is being remachined to accommodate a larger O-ring, in hopes of gaining a solid seal. Thermal model refinements predict excessive solar heating so an MLI (multilayer insulation) heat shield is being installed in front of TIM. Electrical noise and gain testing is in progress and improvements have been made in grounding. The photodiodes were replaced with a superior set. TIM STIM has been tested. The TIM DSP and MU are in initial functional testing, learning curves on the generic channel interface and the digital signal processor have been overcome, and remaining performance glitches are being ironed out. TIM still needs to pass vacuum test, and temperature regulation (PID filter) coefficients for cone and heat sink need to be refined. TIM must undergo environmental testing and thermal vacuum testing with temperature regulation. Post thermal characterizations for noise, standard watt measurement, TIM STIM, and heliostat excitation are planned.

Page 1